Karlottalotta:

Two ropes, AD and BD, are tied to a peg on the ground at point D. The other ends of the ropes are tied to points A and B on a flagpole, as shown below: Two ropes, AD and BD, are tied to a peg on the ground at point D. The other ends of the ropes are tied to points A and B on a f Angle ADC measures 60° and angle BDC measures 45°. What is the distance between the points A and B on the flagpole? (1 point) 6.59 feet 1.73 feet 4.32 feet 2.56 feet I really appreciate all of the help that I have received from you guys

Karlottalotta:

1 attachment
Karlottalotta:

@kacchan

jhonyy9:

sin 60 = AC / DC => DC = AC /(sin 60) => DC = DC => AC/(sin 60) = BC/(sin 45) sin 45 = BC / DC => DC = BC/(sin 45) AC = AB + BC (AB +BC)/(sin 60) = BC/(sin 45) AB +BC = BC*(sin 60)/(sin 45) AB = BC*(sin 60)/(sin 45) - BC

Karlottalotta:

so its c?

jhonyy9:

you need to know the value of sin 60 = sqrt3 /2 and sin 45 = sqrt2 /2

Karlottalotta:

yeah, so c?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 sin 60 = AC / DC => DC = AC /(sin 60) => DC = DC => AC/(sin 60) = BC/(sin 45) sin 45 = BC / DC => DC = BC/(sin 45) AC = AB + BC (AB +BC)/(sin 60) = BC/(sin 45) AB +BC = BC*(sin 60)/(sin 45) AB = [BC*(sin 60)/(sin 45)] - BC \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Karlottalotta yeah, so c? \(\color{#0cbb34}{\text{End of Quote}}\) you need make the calcule - i ve wrote just the way how you get it

jhonyy9:

@Laylalyssa

Karlottalotta:

Im very confused

Karlottalotta:

ahh! its 6.59

Karlottalotta:

Thank you for all of your help!

kacchan:

mhm

Karlottalotta:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kacchan mhm \(\color{#0cbb34}{\text{End of Quote}}\) you just dipped my guy

kacchan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Karlottalotta \(\color{#0cbb34}{\text{Originally Posted by}}\) @kacchan mhm \(\color{#0cbb34}{\text{End of Quote}}\) you just dipped my guy \(\color{#0cbb34}{\text{End of Quote}}\) i said the answer but i deleted it bc i got reported again

kacchan:

and i just got reported as a sent that to you oml

Karlottalotta:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kacchan and i just got reported as a sent that to you oml \(\color{#0cbb34}{\text{End of Quote}}\) ooooo they bogus for that

kacchan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Karlottalotta \(\color{#0cbb34}{\text{Originally Posted by}}\) @kacchan and i just got reported as a sent that to you oml \(\color{#0cbb34}{\text{End of Quote}}\) ooooo they bogus for that \(\color{#0cbb34}{\text{End of Quote}}\) its bc i gave a direct answer when im only supposed to help you not give you the answer

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Karlottalotta Thank you for all of your help! \(\color{#0cbb34}{\text{End of Quote}}\) np anytime --- yw